OpenStudy (anonymous):

What is the height of a rectangular solid with the following dimensions: Length= 7 feet Width= 5 feet Volume= 105 cubic feet

OpenStudy (ganpat):

V = LBH.. So, 105 = 7 * 5 * H Therefore H = 7 ft...

OpenStudy (ganpat):

I have a hint for you.. always look at units first.. they might help you to get the required formula.. over here V= Ft3.. so on the other side it should be ft *ft* ft or ft * ft2.. Hope u get this..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

But @Ganpat if there is a constant value then the answer may be wrong..

OpenStudy (anonymous):

Height is not coming out to 7 feet...

OpenStudy (ganpat):

@waterineyes :yes.. for constant there may be issues, this is just to get tentative idea on what u r dealing with.. and you have take width 7ft.. its actually .. 5ft in question..

OpenStudy (anonymous):

Sorry but I have put it right in the formula...

OpenStudy (anonymous):

@Ganpat your answer is wrong for height i think...

OpenStudy (anonymous):

\[V=Length∗Width∗Height\] Length = 7 Width = 5 V = 105 So, \[105=7×5×H\] \[105=35×H\] \[H =\frac{\cancel{105}^3}{\cancel{35}^1} = 3\] Therefore, Height = 3 Feet...

OpenStudy (ganpat):

oh sorry.. Visual calculations, lil error.. sorry for that.. @xLynnBbyx: Keep a update of this one, my answer was wrong in calculation..